OpenStudy (anonymous):

the 1/2 life of U-235 is 7 x 10^8 years. How much of an original 25.0g sample of U-235 would after 2.1 x 10^9 years ?

OpenStudy (anonymous):

devide 2.1*10^9 by 1/2life to get how many half lifes the sample has undergone then use this equation W=w(1/2)^n where n is the number of halflives W=the weight at the final w= weight at the begining \\ or u could just try it steo by step

OpenStudy (anonymous):

huh ? can u solved for me please

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

n= 2.1*10^9 / 7*10^8 n=3 W = 25g 8(1/2)^3 W = 25/8 g

OpenStudy (compphysgeek):

alternatively, if you know the halflife of Uran, you can use the information to solve the equation \[1/2 w_0 = w_0 \exp(-k T_{1/2}) \] for k. once you've got k you can use the equation again to calculate the weight after \(t = 2.1 \times 10^9\)

OpenStudy (anonymous):

thanks guy

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

if only 0.050mg of PU-239 remains after 6 half lives have past ,how much did the original sample weigh ?

OpenStudy (anonymous):

can u guys help me this one too any idea?

OpenStudy (compphysgeek):

double the amount 6 times ;) the idea is if there is only half of a substance left after halflife, then there was twice as much there at the beginning.

OpenStudy (anonymous):

2 (0.050)^6 ?

OpenStudy (compphysgeek):

after 6 halflifes you've got 0.05, so you've got 0.1 after 5 hl. 0.2 after 4, 0.4 after 3, 0.8 after 2, 1.6 after 1 and 3.2 at the beginning. so it's basically 0.05*2^6 if you use for formular the 0.05^6 gets very small.